Random samples with size 5 are drawn from the population containing the values 26, 32, 41, 50, 58, and 63.


Determine the standard error of the sample means

1
Expert's answer
2022-05-04T14:33:03-0400

We have population values 26, 32, 41, 50, 58, and 63 population size N=6 and sample size n=5.

Mean of population "(\mu)" = 

"\dfrac{26+32+41+50+58+63}{6}=45"


Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{N}=\dfrac{1064}{6}"


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{1064}{6}}"

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{6}C_5=6."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 26, 32, 41, 50, 58 & 207/5 \\
 \hdashline
 2 & 26, 32, 41, 50, 63 & 212/5 \\
 \hdashline
 3 & 26, 32, 41, 63, 58 & 220/5 \\
 \hdashline
 4 & 26, 32, 63, 50, 58 & 229/5 \\
 \hdashline
 5 & 26, 63, 41, 50, 58 & 238/5 \\
 \hdashline
 6 & 63, 32, 41, 50, 58 & 244/5 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) &\bar{X}^2 f(\bar{X})\\ \hline
 207/5 & 1/6 & 207/30 & 42849/150\\
 \hdashline
 212/5 & 1/6 & 212/30 & 44944/150\\
 \hdashline
 220/5 & 1/6 & 220/30 & 48400/150\\
 \hdashline
 229/5 & 1/6 & 229/30 & 52441/150 \\
 \hdashline
 238/5 & 1/6 & 238/30 & 56644/150\\
 \hdashline
 244/5 & 1/6 & 244/30 & 59536/150 \\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=45=\mu"


The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{304814}{150}-(45)^2=\dfrac{1064}{150}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

The standard error of the sample means is

"\sigma_{\bar{X}}=\sqrt{\dfrac{1064}{150}}\approx2.6633"